# Unraveling additional quantum many-body scars of the spin-$1$ XYXY model with Fock-space cages and commutant algebras

### v1

@@ III.1 Definition and destructive interference mechanism
where $r{=}1,2,{\cdots},L$ (owing to PBC, the largest separation $r_{\rm max}{=}L$ [^1] Note that for r=1r{=}1, the state $|\Omega_{1}\rangle$ coincides with the state $|\mathbb{B}_{1}^{\prime}\rangle$ defined in Eq. (9), however even in that case the tower we will build on top of it in the next section is entirely distinct.. The state $|\Omega_{r}\rangle$ resides in the magnetization sector M=2L+2M{=}{-}2L{+}2 and is built from a superposition of $2L$ distinct configurations $|F_{r}^{i}\rangle$ with spin flips at sites $i$ and i+ri{+}r. All of these $|F_{r}^{i}\rangle$ share the same $\mathcal{\hat{C}}-$parity $({-}1)^{r}$, so the support of $|\Omega_{r}\rangle$ lies entirely on one sublattice of the bipartite Fock-space graph defined by $H_{XY}$. As discussed in Sec. II.2, the hopping action of $H_{XY}$ connects each support node $|F_{r}^{i}\rangle$ to basis states only in the complementary sublattice. However, any such complementary-sublattice node is reached from both $|F_{r}^{i}\rangle$ and $|F_{r}^{i{+}1}\rangle$ with identical matrix-element magnitude but opposite relative sign, inherited from the $({-}1)^{i}$ alternating coefficients in $|\Omega_{r}\rangle$ [see Eq. (10)]. Their amplitude in $H_{XY}|\Omega_{r}\rangle$ therefore cancels exactly, and every would-be leakage path is blocked. Hence $|\Omega_{r}\rangle$ is an exact zero-energy eigenstate of $H_{XY}$. Note that even though this can be shown simply by explicitly analyzing the action of $H_{XY}$ on the states directly without any reference to the Fock-space graph, we have illustrated this in the graph language to highlight the FSC origins of these states. A schematic depiction of this cancellation mechanism is shown in Fig. 3 for $|\Omega_{1}\rangle$ in a $12$-site chain.
@@ III.2.2 Revivals
The state $|r,n\rangle$ has magnetization $M_{r,n}{=}2(n{+}1{-}L)$, making it an eigenstate of the total magnetization. Hence, under the full Hamiltonian $H_{0}$ of Eq. (1), the state $|r,n\rangle$ acquires energy $E_{r,n}{=}2h(n{+}1{-}L)$. Consequently for fixed $r$, the set of states $\{|r,n\rangle\}$ forms an equally spaced tower of $H_{0}$ with spacing $\Delta E_{r,n}{=}2h$. For the special case r=Lr{=}L, the alternating phase structure of Eq. (12) enforces $|L,n\rangle{=}0$ for all odd (even) values of $n$ when $L$ is even (odd), so the effective spacing of that particular tower doubles to $4h$. Note that $|r,n\rangle$ also remains an exact eigenstate of the single-ion anisotropy term $D\sum_{i}(S_{i}^{z})^{2}$ acquiring a uniform energy shift of (L2)D(L{-}2)D that is independent of $n$.
The uniform energy spacing immediately implies perfect dynamical revivals from appropriately chosen initial states. A simple choice for the initial states is $|\psi_{r}^{\text{init}}\rangle{=}\sum_{n}|r,n\rangle$ that are coherent superpositions of states from a single tower. Their explicit form,
@@
reveals that they are inherently low entangled since they are essentially a single quasiparticle on top of a product state; in the large-$L$ limit, their half-chain EE approaches $2\ln(2)$. Under time evolution with $H_{0}$, their fidelity $\mathcal{F}_{r}(t){=}|\langle\psi_{r}^{\text{init}}|e^{-\iota H_{0}t}|\psi_{r}^{\text{init}}\rangle|^{2}$ exhibits perfect revivals with period $T{=}\pi/h$ [or $T{=}\pi/(2h)$ for the r=Lr{=}L case due to the double energy spacing of $4h$]. These persistent oscillations from a simple initial state, as illustrated in Fig. 4, are a definitive dynamical hallmark of scar dynamics. Note that it is in principle also possible to further superpose the states in Eq. (15) for different $r$ and obtain Matrix Product State initial states that exhibit similar revivals, but we will not explicitly illustrate them here.
Figure 4: Fidelity $\mathcal{F}(t){=}|\langle\psi(t)|\psi(0)\rangle|^{2}$ for a spin-11 chain of size 2L=122L{=}12 sites, shown for initial states $|\psi_{r}^{\rm init}\rangle$ [see Eq. (15)] (with $r{=}1,2,{\cdots},6$) and $|\psi_{\rm vol}^{\rm init}\rangle$ [see Eq. (31)] constructed as a coherent superposition over the scar eigenstates of the towers $|r,n\rangle$, and the volume-entangled tower $|\mathbb{V}_{n}\rangle$, respectively, and evolved under $H_{0}$ [see Eq. (1)]. All cases exhibit persistent fidelity revivals with period $\pi/h$, consistent with the uniform energy spacing $2h$, except for $|\psi_{r{=}L}^{\rm init}\rangle$, which oscillates with period $\pi/(2h)$ as discussed in the main text.
Figure 5: Bipartite entanglement entropy (EE) of all scar states constructed in this work for a spin-11 chain of size 2L=122L{=}12 sites. The plot shows EE of the tower states $\{|r,n\rangle\}$ [see Eq. (12)] for $r{=}1,2,{\cdots},6$, and the special volume-entangled tower states $\{|\mathbb{V}_{n}\}\rangle$ [see Eq. (25)], both for the standard bipartition with a cut at $L$, and the fine-tuned antipodal bipartition, discussed in the text. For comparison, the EE of all eigenstates of a non-integrable Hamiltonian $H_{\rm comm}^{(2)}$ [see Eq. (22)] with J=h=1J{=}h{=}1, $\epsilon_{i}{\in}[0,0.2]$ and periodicity r=2r{=}2, is shown in the background as grey dots with color-coded by the density of states with a given $E$ (darker=higher density). The black dotted line marks the Page value, the expected EE of a random state. Although the scar states lie in different magnetization sectors and are exact eigenstates of the spin-$1$ XYXY model with different integrability-breaking perturbations, this plot provides a standard comparative representation of their subthermal $E$, highlighting their nonthermal nature.
@@ III.2.5 Connections to other unified frameworks
# QMBS family
$M$
Minimal set of non-
commuting local operators
Bipartite $E$
characteristic
# 1
Bimagnon tower
$|\mathbb{B}_{n}\rangle$, Eq. (8),
Ref. [32]
$n{=}0,1,{\cdots},2L$ 2(nL)2(n{-}L) $\{h_{i,i{+}1}\}$, Eq. (19)
subextensive
Fig. 2 of Ref. [32]
# 2
Bond-bimagnon tower
$|\mathbb{B}^{\prime}_{n}\rangle$, Eq. (9),
Ref. [32, 33]
$n{=}0,1,{\cdots},2L$ 2(nL)2(n{-}L) -
subextensive
Fig. 5 of Ref. [33]
# 3
FSC towers
$|r,n\rangle$, Eq. (12)
For each $r{\in}\{1,{\cdots},L\},$
$n{=}0,{\cdots},2L{-}2$
2(n+1L)2(n{+}1{-}L)
(i) $2L~{\rm mod}~r{=}0$: $\{\mathcal{H}_{j}^{(r)}\}$,
Eq. (20)
(ii) generic r>1r{>}1: $\{q_{i,i{+}1}\}$,
r=1r{=}1: $\{q_{i,i{+}3}\}$
Eq. (23)
subextensive
Fig. 5
# 4
Volume-entangled
tower $|\mathbb{V}_{n}\rangle$, Eq. (25)
$n{=}0,1,{\cdots},2L$ 2(nL)2(n{-}L) $\{\mathcal{H}_{j}^{(L)}\}$ (20)
(i) standard cut:
extensive, Eqs. (27), (28)
(ii) fine-tuned cut:
subextensive, Eqs. (29), (30)
# 5
Mirror-dimer
states $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$, Eq. (33)
For each $k{\in}\{1,{\cdots},L\},$
$m,m^{\prime}{\in}\{{-}1,0,1\}$
$m{+}m^{\prime}$ $\{\mathcal{M}_{j}^{k}\}$, Eq. (32)
(i) dimers cut: $L\ln(3)$
(ii) mirror-paired: 0
@@ IV.1.2 Illustration with the bimagnon tower of scars
Note that the commutant $\mathscr{C}_{0}$ of this particular algebra $\mathscr{A}_{0}$ contains many more symmetries than just the scar projectors, which lead to more complicated block decompositions of the Hilbert space that obscure the clear separation into “thermal” and “non-thermal” subspaces. Nevertheless, this shows that even though the bimagnon states might not be referred to as scars of $H_{XY}$ since they might be explained by other conventional symmetries of this particular model, many local perturbations can be constructed from the generators of $\mathscr{A}_{0}$ that break the extra symmetries while preserving these states, which makes them scars of the perturbed models. Indeed, previous works studying this model [32, 77] add perturbations such as $\sum_{i}((S_{i}^{z})^{2}S_{i+1}^{z}{+}S_{i}^{z}(S_{i+1}^{z})^{2})$, $\sum_{i}(S_{i}^{z})^{2}$ to make these bimagnon states as scars.
@@ IV.1.3 Beyond the bimagnon towers
In the following subsection, we demonstrate how the simplest such choice—formed by periodically clustering the nearest-neighbor bonds $\{h_{i,i+1}\}$ into groups whose total sum reconstructs $H_{XY}$—naturally accounts for a large subset of the $|r,n\rangle$ introduced earlier. Slight modifications of this idea further yield additional families of exact eigenstates, most notably a volume-entangled tower and a set of mirror-dimer states, demonstrating the versatility of the commutant-inspired approach. We summarize the examples and results in Table 2.
Note that since these generators are no longer strictly local operators or uniform sums of such operators, this generalizes the constructions of bond algebras in [77]. However, these generators are nevertheless sums of such strictly local operators; hence, all statements about them not having a unique local Hamiltonian with them as eigenstates remain valid, which, as we have discussed earlier, is sufficient to qualify these states as scars. We also do not attempt to construct the complete algebra whose commutant contains only the scar projectors, which would be a much more formidable task. Instead, we explicitly build perturbations that break all simple symmetries of $H_{0}$ while leaving these states as exact eigenstates. Given the form of such perturbations, it seems highly unlikely that there is any simple symmetry that would survive and explain the existence of these eigenstates.
@@ IV.2.1 Clustering approach
where $r$ is a non-trivial divisor of the system size $2L$ (i.e. $2{\leq}r{\leq}L$), such that $H_{XY}{=}\sum_{j{=}1}^{r}\mathcal{H}_{j}^{(r)}$. Each $\mathcal{H}_{j}^{(r)}$ is the sum of $2L/r$ non-overlapping exchange terms separated by $r$ sites, and the family $\{\mathcal{H}_{j}^{(r)}\}$ provides a minimal non-commuting generator set whose direct sum yields $H_{XY}$. We then search for simultaneous eigenstates of all $\mathcal{H}_{j}^{(r)}$. Note that we excluded divisors r=1r{=}1 and r=2Lr{=}2L since r=1r{=}1 amounts to considering the entire Hamiltonian as single block while r=2Lr{=}2L involves considering each $h_{i,i{+}1}$, which simply yields the bimagnon towers as we explained previously [see Eqs. (19) and (8)].
Remarkably, for every allowed divisor $r$ we find that the tower states $\{|r,n\rangle\}$ satisfy (see Appendix. C.1)
@@
with arbitrary periodic coefficients $\epsilon_{j}{=}\epsilon_{j{+}r}$. The states $\{|r,n\rangle\}$ still remain exact eigenstates of $H_{\rm comm}^{(r)}{=}H_{0}{+}V_{\text{comm}}^{(r)}$ with energy unchanged at $E_{r,n}{=}2h(n{+}1{-}L)$. Meanwhile, the distribution of adjacent gap ratios $r_{\alpha}{=}\text{min}(\delta_{\alpha{+}1}/\delta_{\alpha},\delta_{\alpha}/\delta_{\alpha{+}1})$ with $\delta_{\alpha}{=}\mathcal{E}_{\alpha}{-}\mathcal{E}_{\alpha{-}1}$, where $\mathcal{E}_{\alpha}\}$ are eigenvalues of $H_{\rm comm}^{r}$, follows the Wigner-Dyson distribution associated with Gaussian Orthogonal Ensemble (GOE), confirming that the perturbed model is nonintegrable [see Fig. 9 (a)]. This same chaotic reference Hamiltonian was used in Fig. 5 for the EE comparison, where the states $|r,n\rangle$ persist as robust low-entropy outliers in an otherwise thermal spectrum.

### v2

@@ III.1 Definition and destructive interference mechanism
where $r{=}1,2,{\cdots},L$ (owing to PBC, the largest separation $r_{\rm max}{=}L$ [^1] Note that for r=1r{=}1, the state $|\Omega_{1}\rangle$ coincides with the state $|\mathbb{B}_{1}^{\prime}\rangle$ defined in Eq. (9), however even in that case the tower we will build on top of it in the next section is entirely distinct.. The state $|\Omega_{r}\rangle$ resides in the magnetization sector M=2L+2M{=}{-}2L{+}2 and is built from a superposition of $2L$ distinct configurations $|F_{r}^{i}\rangle$ with spin raised at sites $i$ and i+ri{+}r. All of these $|F_{r}^{i}\rangle$ share the same $\mathcal{\hat{C}}-$parity $({-}1)^{r}$, so the support of $|\Omega_{r}\rangle$ lies entirely on one sublattice of the bipartite Fock-space graph defined by $H_{XY}$. As discussed in Sec. II.2, the hopping action of $H_{XY}$ connects each support node $|F_{r}^{i}\rangle$ to basis states only in the complementary sublattice. However, any such complementary-sublattice node is reached from both $|F_{r}^{i}\rangle$ and $|F_{r}^{i{+}1}\rangle$ with identical matrix-element magnitude but opposite relative sign, inherited from the $({-}1)^{i}$ alternating coefficients in $|\Omega_{r}\rangle$ [see Eq. (10)]. Their amplitude in $H_{XY}|\Omega_{r}\rangle$ therefore cancels exactly, and every would-be leakage path is blocked. Hence $|\Omega_{r}\rangle$ is an exact zero-energy eigenstate of $H_{XY}$. Note that even though this can be shown simply by explicitly analyzing the action of $H_{XY}$ on the states directly without any reference to the Fock-space graph, we have illustrated this in the graph language to highlight the FSC origins of these states. A schematic depiction of this cancellation mechanism is shown in Fig. 3 for $|\Omega_{1}\rangle$ in a $12$-site chain.
@@ III.2.2 Revivals
The state $|r,n\rangle$ is an eigenstate of the total magnetization with magnetization $M_{r,n}{=}2(n{+}1{-}L)$. Hence, under the full Hamiltonian $H_{0}$ of Eq. (1), the state $|r,n\rangle$ acquires energy $E_{r,n}{=}2h(n{+}1{-}L)$. Consequently for fixed $r$, the set of states $\{|r,n\rangle\}$ forms an equally spaced tower of $H_{0}$ with spacing $\Delta E_{r,n}{=}2h$. For the special case r=Lr{=}L, the alternating phase structure of Eq. (12) enforces $|L,n\rangle{=}0$ for all odd (even) values of $n$ when $L$ is even (odd), so the effective spacing of that particular tower doubles to $4h$. Note that $|r,n\rangle$ also remains an exact eigenstate of the single-ion anisotropy term $D\sum_{i=1}^{2L}(S_{i}^{z})^{2}$ under which it acquires a uniform energy shift of (L2)D(L{-}2)D that is independent of $n$.
The uniform energy spacing immediately implies perfect dynamical revivals from appropriately chosen initial states. A simple choice for the initial states is $|\psi_{r}^{\text{init}}\rangle{=}\sum_{n=0}^{2L-2}|r,n\rangle$ that that are coherent superpositions of states from a single tower. Their explicit form,
@@
reveals that they are inherently low entangled since they are essentially a single quasiparticle on top of a product state; in the large-$L$ limit, their half-chain EE approaches $2\ln(2)$. Under time evolution with $H_{0}$, their fidelity $\mathcal{F}_{r}(t){=}|\langle\psi_{r}^{\text{init}}|e^{-\iota H_{0}t}|\psi_{r}^{\text{init}}\rangle|^{2}$ exhibits perfect revivals with period $T{=}\pi/h$ [or $T{=}\pi/(2h)$ for the r=Lr{=}L case due to the double energy spacing of $4h$]. These persistent oscillations from a simple initial state, as illustrated in Fig. 4, are a definitive dynamical hallmark of scarring. Note that it is in principle also possible to further superpose the states in Eq. (15) for different $r$ and obtain Matrix Product State initial states that exhibit similar revivals, but we will not explicitly illustrate them here.
Figure 4: Fidelity $\mathcal{F}(t){=}|\langle\psi(t)|\psi(0)\rangle|^{2}$ for a spin-11 chain of size 2L=122L{=}12 sites, shown for initial states $|\psi_{r}^{\rm init}\rangle$ [see Eq. (15)] (with $r{=}1,2,{\cdots},6$) and $|\psi_{\rm vol}^{\rm init}\rangle$ [see Eq. (32)] constructed as a coherent superposition over the scar eigenstates of the towers $|r,n\rangle$, and the volume-entangled tower $|\mathbb{V}_{n}\rangle$, respectively. The state $|\psi_{r}^{\rm init}\rangle$ is time-evolved with $H_{0}$ [see Eq. (1)], while $|\psi_{\rm vol}^{\rm init}\rangle$ is time-evolved with $H_{\rm vol}$ [see Eq. (25)]. All cases exhibit persistent fidelity revivals with period $\pi/h$, consistent with the uniform energy spacing $2h$, except for $|\psi_{r{=}L}^{\rm init}\rangle$, which oscillates with period $\pi/(2h)$ as discussed in the main text.
Figure 5: Bipartite entanglement entropy (EE) of all scar states constructed in this work for a spin-11 chain of size 2L=122L{=}12 sites. The plot shows EE of the tower states $\{|r,n\rangle\}$ [see Eq. (12)] for $r{=}1,2,{\cdots},6$, and the special volume-entangled tower states $\{|\mathbb{V}_{n}\rangle\}$ [see Eq. (25)], both for the standard bipartition with a cut at $L$, and the fine-tuned antipodal bipartition, discussed in the text. For comparison, the EE of all eigenstates of a non-integrable Hamiltonian $H_{\rm comm}^{(2)}$ [see Eq. (22)] with J=h=1J{=}h{=}1, $\epsilon_{i}{\in}[0,0.2]$ and periodicity r=2r{=}2, is shown in the background as grey dots with its color encoding by the density of states with a given $E$ (darker=higher density). The black dotted line marks the Page value, the expected EE of a random state. Although the scar states lie in different magnetization sectors and are exact eigenstates of the spin-$1$ XYXY model with different integrability-breaking perturbations, this plot provides a standard comparative representation of their subthermal $E$, highlighting their nonthermal nature.
@@ III.2.5 Connections to other unified frameworks
# QMBS family $M$ Minimal set of non- commuting local operators Bipartite EE characteristic
# 1 Bimagnon tower $|\mathbb{B}_{n}\rangle$, Eq. (8), Ref. [31] $n{=}0,1,{\cdots},2L$ 2(nL)2(n{-}L) $\{h_{i,i{+}1}\}$, Eq. (19) subextensive Fig. 2 of Ref. [31]
# 2 Bond-bimagnon tower $|\mathbb{B}^{\prime}_{n}\rangle$, Eq. (9), Ref. [31, 32] $n{=}0,1,{\cdots},2L$ 2(nL)2(n{-}L) - subextensive Fig. 5 of Ref. [32]
# 3 FSC towers $|r,n\rangle$, Eq. (12) For each $r{\in}\{1,{\cdots},L\},$ $n{=}0,{\cdots},2L{-}2$ 2(n+1L)2(n{+}1{-}L) (i) $2L~{\rm mod}~r{=}0$: $\{\mathcal{H}_{j}^{(r)}\}$, Eq. (20) (ii) generic r>1r{>}1: $\{q_{i,i{+}1}\}$, r=1r{=}1: $\{q_{i,i{+}3}\}$ Eq. (23) subextensive Fig. 5
# 4 Volume-entangled tower $|\mathbb{V}_{n}\rangle$, Eq. (25) $n{=}0,1,{\cdots},2L$ 2(nL)2(n{-}L) $\{\mathcal{H}_{j}^{(L)}\}$ (20) (i) standard cut: extensive, Eqs. (28), (29) (ii) fine-tuned cut: subextensive, Eqs. (30), (31)
# 5 Mirror-dimer states $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$, Eq. (34) For each $k{\in}\{1,{\cdots},L\},$ $^{\prime}{\in}\{{-}1,0,1\} m+mm{+}m^{\prime}$ $\{\mathcal{M}_{j}^{k}\}$, Eq. (33) (i) dimers cut: $L\ln(3)$ (ii) mirror-paired: 0
@@ IV.1.2 Illustration with the bimagnon tower of scars
Note that the commutant $\mathscr{C}_{0}$ of this particular algebra $\mathscr{A}_{0}$ contains many more symmetries than just the scar projectors, which lead to more complicated block decompositions of the Hilbert space that obscure the clear separation into “thermal” and “non-thermal” subspaces. Nevertheless, this shows that even though the bimagnon states might not be referred to as scars of $H_{XY}$ since they might be explained by other conventional symmetries of this particular model, many local perturbations can be constructed from the generators of $\mathscr{A}_{0}$ that break the extra symmetries while preserving these states, which makes them scars of the perturbed models. Indeed, previous works studying this model [31, 74] add perturbations such as $\sum_{i=1}^{2L}((S_{i}^{z})^{2}S_{i+1}^{z}{+}S_{i}^{z}(S_{i+1}^{z})^{2})$, $\sum_{i=1}^{2L}(S_{i}^{z})^{2}$ to make these bimagnon states as scars.
@@ IV.1.3 Beyond the bimagnon towers
In the following subsection, we demonstrate how the simplest such choice—formed by periodically clustering the nearest-neighbor bonds $\{h_{i,i+1}\}$ into groups whose total sum reconstructs $H_{XY}$—naturally accounts for a large subset of the $|r,n\rangle$ states introduced earlier. Slight modifications of this idea further yield additional families of exact eigenstates, most notably a volume-entangled tower and a set of mirror-dimer states, demonstrating the versatility of the commutant-inspired approach. We summarize the examples and results in Table 2.
Note that since these generators are no longer strictly local operators or uniform sums of such operators, this generalizes the constructions of bond algebras of Ref. [74]. However, these generators are nevertheless sums of such strictly local operators; hence, all statements about their singlets not having a unique local Hamiltonian with the singlets as eigenstates remain valid, which, as we have discussed earlier, is sufficient to qualify these singlets as scars. We also do not attempt to construct the complete algebra whose commutant contains only the scar projectors, which would be a much more formidable task. Instead, we explicitly build perturbations that break all simple symmetries of $H_{0}$ while leaving these states as exact eigenstates. Given the form of such perturbations, it seems highly unlikely that there is any simple symmetry that would survive and explain the existence of these eigenstates.
@@ IV.2.1 Clustering approach
where $r$ is a non-trivial divisor of the system size $2L$ (i.e. $2{\leq}r{\leq}L$), such that $H_{XY}{=}\sum_{j{=}1}^{r}\mathcal{H}_{j}^{(r)}$. Each $\mathcal{H}_{j}^{(r)}$ is the sum of $2L/r$ non-overlapping exchange terms separated by $r$ sites, and the family of operators $\{\mathcal{H}_{j}^{(r)}\}$ provides a minimal non-commuting generator set whose direct sum yields $H_{XY}$. We then search for simultaneous eigenstates of all $\mathcal{H}_{j}^{(r)}$. Note that we excluded divisors r=1r{=}1 and r=2Lr{=}2L since r=1r{=}1 amounts to considering the entire Hamiltonian as single block while r=2Lr{=}2L involves considering each $h_{i,i{+}1}$, which simply yields the bimagnon towers as we explained previously [see Eqs. (19) and (8)].
Remarkably, for every allowed divisor $r$ we find that the tower states $\{|r,n\rangle\}$ of Eq. (12) satisfy (see Appendix. C.1)
@@
with arbitrary periodic coefficients $\epsilon_{j}{=}\epsilon_{j{+}r}$. The states $\{|r,n\rangle\}$ still remain exact eigenstates of $H_{\rm comm}^{(r)}{=}H_{0}{+}V_{\text{comm}}^{(r)}$ with energy unchanged at $E_{r,n}{=}2h(n{+}1{-}L)$. Meanwhile, the distribution of adjacent gap ratios $r_{\alpha}{=}\text{min}(\delta_{\alpha{+}1}/\delta_{\alpha},\delta_{\alpha}/\delta_{\alpha{+}1})$ with $\delta_{\alpha}{=}\mathcal{E}_{\alpha}{-}\mathcal{E}_{\alpha{-}1}$, where $\mathcal{E}_{\alpha}\}$ are eigenvalues of $H_{\rm comm}^{(r)}$, follows the Wigner-Dyson distribution associated with Gaussian Orthogonal Ensemble of Random Matrix Theory, confirming that the perturbed model is nonintegrable [see Fig. 9 (a)]. This same chaotic reference Hamiltonian was used in Fig. 5 for the EE comparison, where the states $|r,n\rangle$ persist as robust low-entropy outliers in an otherwise thermal spectrum.